OpenStudy (anonymous):

Which values of x will make the absolute value equation true? |x – 3| = 40 A. {–43, –37} B. {43, –37} C. {40, –40} D. {43, 37}

OpenStudy (anonymous):

You do this problem in two parts. For the equation to be true, the argument of the absolute value (the expression inside) has to be either forty, or the opposite of forty. I.e.,\[|x – 3| = 40 \implies x-3=40~or~x-3=-40\] Solve those two equations to get your answer.